She spends her entire monthly food allowance Rs.42 on rice and wheat flour. The price of rice per kg RS.2 and the price of wheat flour is Re.1 per kg. She purchase 12 kg of rice and 18 kg of wheat flour and her marginal rate of substitution between rice and wheat flour is Rs.1 . Is she in equilibrium? Explain.
1
Expert's answer
2020-06-29T14:44:51-0400

Let "I=42, p_x=2,p_y=1,Q_x=12,Q_y=18"

"MRS_{xy}=\frac {MU_x}{MU_y}=1"

"MU_x=MU_y"


"p_xQ_x+p_yQ_y=I"

Let us verify the fulfillment of this condition and the conditions for maximizing utility.


"2\times12+1\times18=42"


"\frac {MU_x}{p_x}=\frac {MU_y}{p_y}"


"\frac {MU_x}{2}{=}\mathllap{/\,} \frac{MU_y}{1}"

She is not in balance